A nitrogen sample has a volume of 6 litres at 35 degrees celsius and a pressure of 0,974.What volume will it occupy at STP


1
Expert's answer
2022-04-29T04:10:01-0400

"(P1V1)\/T1 = (P2V2)\/T2,, (0.974*6)\/308.15 = (V(str)*1)\/298.15, V(str) =5.65 liters"